OpenStudy (anonymous):

The line x+2y=9 intersects the curve xy +18=0 at the points A and B. Find the coordinates of A and B.

sam (.sam.):

Simultaneous

sam (.sam.):

 you can see both equations intercept each other, this means they will have the same x and y values, so x+2y=9 x=9-2y and xy +18=0 x=-18/y Then, equate both together 9-2y=-18/y Then solve